# Entropy-Regularized Token-Level Policy Optimization for Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have shown promise as intelligent agents in interactive decision-making tasks. Recent approaches often depend on meticulously designed prompts, high-quality examples, or additional reward models for in-context learning, supervised fine-tuning, or RLHF. Reinforcement learning (RL) presents a dynamic alternative for LLMs to overcome these dependencies by engaging directly with task-specific environments. Nonetheless, it faces significant hurdles: 1) instability stemming from the exponentially vast action space requiring exploration; 2) challenges in assigning token-level credit based on action-level reward signals, resulting in discord between maximizing rewards and accurately modeling corpus data. In response to these challenges, we introduce Entropy-Regularized Token-level Policy Optimization (ETPO), an entropy-augmented RL method tailored for optimizing LLMs at the token level. At the heart of ETPO is our novel per-token soft Bellman update, harmonizing RL process with the principles of language modeling. It decomposes the Q-function update from a coarse action-level view to a more granular token-level perspective, backed by theoretical proof of optimization consistency. Crucially, this decomposition renders linear time complexity in action exploration. We assess the effectiveness of ETPO within a simulated environment that models data science code generation as a series of multi-step interactive tasks; results show that ETPO achieves effective performance improvement on the CodeLlama-7B model and surpasses a variant PPO baseline inherited from RLHF. This underlines ETPO’s potential as a robust method for refining the interactive decision-making capabilities of LLMs. Our code is open-sourced at https://github.com/morning9393/ETPO.
@@ 4 Entropy-Regularized Token-Level Policy Optimization
$Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})=\begin{cases}\mathbb{E}_{w_{t}^{j+1}\sim\pi(s_{t},w_{t}^{1:j})}[Q(s_{t},w_{t}^{1:j},w_{t}^{j+1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t},w_{t}^{1:j}),&\mbox{if }j<|a_{t}|\\ r(s_{t},a_{t})+\gamma(\mathbb{E}_{w_{t+1}^{1}\sim\pi(s_{t+1})}[Q(s_{t+1},w_{t+1}^{1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t+1})),&\mbox{if }j=|a_{t}|\end{cases}$ (6)
$\pi_{new}=\arg\min_{\pi}\mathcal{J}_{Q}(\pi)=\arg\min_{\pi}D_{KL}\Big{[}\pi(\cdot|s_{t},w_{t}^{1:j-1})||\exp(Q(s_{t},w_{t}^{1:j-1},\cdot)/\beta)\Big{]}.$ (7)
@@ 4.2 Per-Token Policy Update
Recall the Equation 13, we rewrite this optimal policy at the token level:
$\pi^{*}(\cdot|s_{t},w_{t}^{1:j-1})=\frac{\bar{\pi}(w_{t}^{j}|s_{t},w_{t}^{1:j-1})\exp(Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})/\beta)}{\mathbb{E}_{\bar{w}_{t}^{j}\sim\bar{\pi}}[\exp(Q(s_{t},w_{t}^{1:j-1},\bar{w}_{t}^{j})/\beta)]}.$ (9)
Since $\bar{\pi}(w_{t}^{j}|s_{t},w_{t}^{1:j-1})/\mathbb{E}_{\bar{w}_{t}^{j}\sim\bar{\pi}}[\exp(Q(s_{t},w_{t}^{1:j-1},\bar{w}_{t}^{j})/\beta)]$ is a normalizing constant that does not contribute to the gradient concerning the new policy and can thus be ignored (Haarnoja et al., 2018b), we have $\pi^{*}(w_{t}^{j}|s_{t},w_{t}^{1:j-1})\propto\exp(Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})/\beta)$ and can thus update the policy toward minimizing the KL-divergence between $\pi(w_{t}^{j}|s_{t},w_{t}^{1:j-1})$ and $\exp(Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})/\beta)$, which is somewhat similar to the policy update process in SAC. Our per-token policy update scheme is to minimize the objective $\mathcal{J}_{Q}(\pi)$ in Equation 7.

### v4

# Entropy-Regularized Token-Level Policy Optimization for Language Agent Reinforcement

@@ Abstract
Large Language Models (LLMs) have shown promise as intelligent agents in interactive decision-making tasks. Recent approaches often depend on meticulously designed prompts, high-quality examples, or additional reward models for in-context learning, supervised fine-tuning, or RLHF. Reinforcement learning (RL) presents a dynamic alternative for LLMs to overcome these dependencies by engaging directly with task-specific environments. Nonetheless, it faces significant hurdles: 1) instability stemming from the exponentially vast action space requiring exploration; 2) challenges in assigning token-level credit based on action-level reward signals, resulting in discord between maximizing rewards and accurately modeling corpus data. In response to these challenges, we introduce Entropy-Regularized Token-level Policy Optimization (ETPO), an entropy-augmented RL method tailored for optimizing LLMs at the token level. At the heart of ETPO is our novel per-token soft Bellman update, harmonizing RL process with the principles of language modeling. It decomposes the Q-function update from a coarse action-level view to a more granular token-level perspective, backed by theoretical proof of optimization consistency. Crucially, this decomposition renders linear time complexity in action exploration. We assess the effectiveness of ETPO within a simulated environment that models data science code generation as a series of multi-step interactive tasks; results show that ETPO achieves effective performance improvement on the CodeLlama-7B model and surpasses a variant PPO baseline inherited from RLHF. This underlines ETPO’s potential as a robust method for refining the interactive decision-making capabilities of LLMs [^1] Our code is open-sourced at https://github.com/morning9393/ETPO. For a more detailed preliminary work describing our motivation for token-level decomposition and applying it in PPO methods, please refer to Wen et al. (2024).
@@ 4 Entropy-Regularized Token-Level Policy Optimization
$Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})\leftarrow\begin{cases}\mathbb{E}_{w_{t}^{j+1}\sim\pi(s_{t},w_{t}^{1:j})}[Q(s_{t},w_{t}^{1:j},w_{t}^{j+1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t},w_{t}^{1:j}),&\mbox{if }j<|a_{t}|\\ r(s_{t},a_{t})+\gamma(\mathbb{E}_{w_{t+1}^{1}\sim\pi(s_{t+1})}[Q(s_{t+1},w_{t+1}^{1})]-\beta D_{KL}[\pi||\bar{\pi}](s_{t+1})),&\mbox{if }j=|a_{t}|\end{cases}$ (6)
$\pi_{new}=\arg\min_{\pi}\mathcal{J}_{Q}(\pi)=\arg\min_{\pi}D_{KL}\Big{[}\pi(\cdot|s_{t},w_{t}^{1:j-1})||\pi_{Q}^{*}(\cdot|s_{t},w_{t}^{1:j-1})\Big{]}.$ (7)
@@ 4.2 Per-Token Policy Update
Recall the Equation 13, we rewrite this optimal policy regarding to the Q function at the token level:
$\pi_{Q}^{*}(\cdot|s_{t},w_{t}^{1:j-1})=\frac{\bar{\pi}(w_{t}^{j}|s_{t},w_{t}^{1:j-1})\exp(Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})/\beta)}{\mathbb{E}_{\bar{w}_{t}^{j}\sim\bar{\pi}}[\exp(Q(s_{t},w_{t}^{1:j-1},\bar{w}_{t}^{j})/\beta)]}.$ (9)
Due to the discrete nature of vocabularies, we can explicitly calculate the $\pi_{Q}^{*}(\cdot|s_{t},w_{t}^{1:j-1})$ with $\bar{\pi}(w_{t}^{j}|s_{t},w_{t}^{1:j-1})$ and current Q function $Q(s_{t},w_{t}^{1:j-1},w_{t}^{j})$, and thus update the policy toward minimizing the KL-divergence between $\pi(w_{t}^{j}|s_{t},w_{t}^{1:j-1})$ and $\pi_{Q}^{*}(\cdot|s_{t},w_{t}^{1:j-1})$, which is somewhat similar to the policy update process in SAC (Haarnoja et al., 2018b). Our per-token policy update scheme is to minimize the objective $\mathcal{J}_{Q}(\pi)$ in Equation 7.